# Why $w\neq-1$? Anthropic Selection in a $\Lambda$ + Axion Dark Energy Model

### v1

@@ III Discussion and Conclusions
So far, we have focused on the case where the bare cosmological constant is restricted to be negative. This simplifying assumption is well-motivated in the context of string theory, where vacua with negative vacuum energy are favored. Nonetheless, it is natural to consider generalizations in which the bare cosmological constant can also take positive values. We have also assumed that observers exist in the universe and observe cosmic acceleration. Relaxing this assumption is another possible extension. If one allows a bare positive cosmological constant or does not impose the condition that observers see an accelerating universe, the anthropic argument no longer provides a lower bound on $m$ when assuming a flat measure in $\ln m$ for the axion mass. This is because the probability distribution remains constant for arbitrarily small $m$, leading to a divergent probability. This requires some form of regularization, either through a modified measure or by imposing a prior on the axion mass.
@@
It should also be noted that including both eternally expanding universes and those that eventually recollapse in the anthropic argument is itself nontrivial. In the present work, we imposed a lower bound on the recollapse time, assuming that the emergence of intelligent observers like ourselves requires a timescale after galaxy formation, comparable to the current age of our universe. Unlike the case of a positive cosmological constant, where galaxies can persist indefinitely once formed, a universe with negative vacuum energy inevitably undergoes recollapse within a finite time. Whether such recollapsing universes should be assigned the same anthropic weight as eternally expanding ones is unclear, and addressing this issue may require a more refined treatment of the multiverse measure. A proper assessment of these issues would involve a careful reevaluation of the prior distributions over both the cosmological constant and the axion parameters. We leave such investigations for future work.
@@
In this Letter, we have explored a dark energy model composed of a bare negative cosmological constant and a single ultra-light axion, motivated by the string axiverse. Assuming that intelligent observers exist and observe an accelerating universe, a conditional anthropic bound (6) naturally follows, which constrains the initial total dark energy density $\rho_{\mathrm{DE}}$ to lie within a range compatible with both galaxy formation and cosmic acceleration. Under this condition, we derived nontrivial anthropic constraints on both the axion mass and the bare cosmological constant. The axion mass is bounded from above to avoid fine-tuning of the initial misalignment angle near the hilltop, and from below to prevent the total dark energy from requiring an unnaturally small bare cosmological constant.
@@
We have found that, within the anthropically allowed region, the axion mass typically lies around ${\cal O}(10)H_{0}$ for a decay constant near the Planck scale, and it increases for smaller decay constants. While the most likely value of the dark energy equation-of-state parameter $w_{0}$ lies close to −$-1$, the deviation by ${\cal O}(0.1)$ is statistically significant. This provides a natural anthropic explanation for why $w_{0}\neq-1$ may be expected, and is intriguingly consistent with recent DESI results hinting at time-varying dark energy Adame et al. (2025); Abdul Karim et al. (2025a, b). Our results thus offer a concrete realization of how the string theory landscape, together with a light axion, can shape the nature of dark energy in our universe.

### v2

@@ III Discussion and Conclusions
So far, we have focused on the case where the bare cosmological constant is restricted to be negative. This simplifying assumption is well-motivated in the context of string theory, where vacua with negative vacuum energy are favored. Nonetheless, it is natural to consider generalizations in which the bare cosmological constant can also take positive values. We have also assumed that observers exist in the universe and observe cosmic acceleration. Relaxing this assumption is another possible extension. If one allows an unrestricted range of positive values for the bare cosmological constant, or does not impose the condition that observers see an accelerating universe, the anthropic argument no longer provides a lower bound on $m$ when assuming a flat measure in $\ln m$ for the axion mass. This is because the probability distribution remains constant for arbitrarily small $m$, leading to a divergent probability. This requires some form of regularization, either through a modified measure or by imposing a prior on the axion mass. However, if the allowed positive values of the cosmological constant are limited to those smaller than, for instance, $\rho_{m,0}/2$, our analysis can be straightforwardly extended without encountering this issue. In such cases, the lower bound on $m$ still emerges from the conditional anthropic condition.
Alternatively, if we restrict ourselves to the case of a vanishing cosmological constant (i.e., the pure axion dark energy model), we obtain qualitatively similar results for the likelihood functions of the dark energy equation-of-state parameter, axion mass, and dark energy density.
@@
It should also be noted that including both eternally expanding universes and those that eventually recollapse in the anthropic argument is itself nontrivial. Unlike the case of a positive cosmological constant, where galaxies can persist indefinitely once formed, a universe with negative vacuum energy inevitably undergoes recollapse within a finite time. Whether such recollapsing universes should be assigned the same anthropic weight as eternally expanding ones is unclear, and addressing this issue may require a more refined treatment of the multiverse measure. A proper assessment of these issues would involve a careful reevaluation of the prior distributions over both the cosmological constant and the axion parameters. We leave such investigations for future work.
We have focused on a dark energy model composed of a cosmological constant and an axion. Axions are also known to play other roles in cosmology, such as contributing to dark matter and dark radiation, and anthropic arguments concerning their abundances have been discussed in Refs. Hellerman and Walcher (2005); Takahashi and Yamada (2019). It would be interesting to extend our analysis to include these additional dark components.
@@
In this paper, we have explored a dark energy model composed of a bare negative cosmological constant and a single ultra-light axion, motivated by the string axiverse. Assuming that intelligent observers exist and observe an accelerating universe, a conditional anthropic bound (7) naturally follows, which constrains the initial total dark energy density $\rho_{\mathrm{DE,i}}$ to lie within a range compatible with both galaxy formation and cosmic acceleration. Under this condition, we derived nontrivial anthropic constraints on both the axion mass and the bare cosmological constant. The axion mass is bounded from above to avoid fine-tuning of the initial misalignment angle near the hilltop, and from below to achieve accelerated expansion.
@@
We have found that, within the anthropically allowed region, the axion mass typically lies around ${\cal O}(10)H_{0}$ for a decay constant near the Planck scale, and it slightly decreases for smaller decay constants (see Appendix A). While the most likely value of the dark energy equation-of-state parameter $w_{0}$ lies close to −$-1$, the deviation by ${\cal O}(0.1)$ is statistically significant. This provides a natural anthropic explanation for why $w_{0}\neq-1$ may be expected, and is intriguingly consistent with recent DESI results hinting at time-varying dark energy Adame et al. (2025); Abdul Karim et al. (2025a, b). Our results thus offer a concrete realization of how the string theory landscape, together with a light axion, can shape the nature of dark energy in our universe.
@@ Acknowledgments
Appendix A Dependence on the parameters
Here, we discuss the dependence of our results on the parameters such as $\rho_{\Lambda}^{\mathrm{max}}$ and $f$.
A.1 Smaller $\rho_{\Lambda}^{\mathrm{max}}$
As discussed in the main text, the mass likelihood for $\ln m$ has a peak at $m=m_{\mathrm{peak}}$ determined by $2m_{\mathrm{peak}}^{2}f^{2}=\rho_{\Lambda}^{\mathrm{max}}$. Thus, if we could set a more stringent upper bound on $\rho_{\Lambda}$ from refined anthropic considerations or other reasoning, $m_{\mathrm{peak}}$ would become smaller. To illustrate this expectation, we show the mass likelihood assuming a test value of $\rho_{\Lambda}^{\mathrm{max}}=2\rho_{\mathrm{DE,obs}}$ in Fig. 5. As expected, we obtain a peak at $m_{\mathrm{peak}}\simeq 1.4H_{0}$.
Figure 5: Same as Fig. 2 except for $\rho_{\Lambda}^{\mathrm{max}}=2\rho_{\mathrm{DE,obs}}$.
A.2 Smaller $f$
Next, we consider smaller decay constants. We show the allowed parameter regions for $f=0.1M_{\mathrm{Pl}}$ with $m=10H_{0}$, $7H_{0}$, and $5H_{0}$ in Fig. 6. For $m=5H_{0}$, $\rho_{\Lambda}$ should be near zero to realize the accelerated universe with the limited energy budget of the axion, $\rho_{\phi}\leq 2m^{2}f^{2}$. Similarly to the bottom panel in Fig. 1, the allowed value of $w_{0}$ does not deviate largely from −$-1$. For $m=7H_{0}$ and $10H_{0}$, larger $w_{0}$ is allowed for large $\rho_{\Lambda}$.
Figure 6: Same as Fig. 1 except for the values of $m$ and $f=0.1M_{\mathrm{Pl}}$.
We show the likelihood for $\ln m$ with $f=0.1M_{\mathrm{Pl}}$ and $0.2M_{\mathrm{Pl}}$ in Fig. 7. The mechanisms of the suppression for small and large masses are the same as in Fig. 2. However, the suppression for large mass is more severe for smaller $f$. This is because the smaller Hubble parameter for the same $m$ requires finer tuning for $\theta_{\mathrm{i}}$. As a result, the peak mass becomes smaller than the estimation by $2m_{\mathrm{peak}}^{2}f^{2}=\rho_{\Lambda}^{\mathrm{max}}$.
Figure 7: Same as Fig. 2 except for $f=0.1M_{\mathrm{Pl}}$ (blue) and $0.2M_{\mathrm{Pl}}$ (red).
Finally, we show the probability distribution for $\rho_{\mathrm{DE,i}}$ with $f=0.1M_{\mathrm{Pl}}$ and $0.2M_{\mathrm{Pl}}$. As mentioned above, the required tuning for $\theta_{\mathrm{i}}$ becomes more severe for smaller $f$. As a result, $\rho_{\mathrm{DE,i}}$ becomes the order of the upper bound $\rho_{\Lambda}^{\mathrm{max}}$ only with a negligibly small probability. Rather, $\rho_{\mathrm{DE,i}}$ typically lies around the observed value $\rho_{\mathrm{DE,obs}}$, which is the minimum order to realize the accelerated expansion. While the probability distribution is concentrated below $\rho_{\rm DE,i}=\rho_{\rm DE,obs}$ for $f=0.1M_{\mathrm{Pl}}$, it spreads around $\rho_{\rm DE,i}\sim\rho_{\rm DE,obs}$ for $f=0.2M_{\mathrm{Pl}}$.
Figure 8: Same as Fig. 4 except for $f=0.1M_{\mathrm{Pl}}$ (blue) and $0.2M_{\mathrm{Pl}}$ (red).